Ellen invested $9000, part at three percent per year and the rest at 5percent per year. After one year the total interest from these investments was $354. How much did she invest at each rate?
I have been looking at this for over an hour andnim not sure even how to start it.. Thank you so much!!

Answer by rfer(16322)

0.03x+0.05(9000-x)=354
0.03x+450-0.05x=354
-0.02x=-96
x=4800 at 3%
9000-x=4200 at 5%